OpenStudy (anonymous):

A car of mass 1200 kg traveling westward at 30. m/s is slowed to a stop in a distance of 50. m by the car’s brakes. What is the acceleration of the car? acceleration = (final velocity squared - initial velocity squared) / (2 times the displacement)

OpenStudy (anonymous):

I don't understand how i find the displacement and isnt the final and initial velocity 30 and 50?

OpenStudy (mayankdevnani):

use 3 equation of motion :- \[\large \bf v^2=u^2+2as\] where, v=0, u=30 m/s s=50 m a=? (you have to calculate) Plug all the values in this equation and get your answer :)

OpenStudy (mayankdevnani):

hope you understand. @dar2015